# Parallel Batch Scheduling With Incompatible Job Families Via Constraint Programming

### v2

@@ VI-C The Redundant Synchronized Model
$x_{j}$$\in\{[a,a+p_{j}:r_{j}\leq a\in\mathbb{Z})\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J};$ $(33)$
$x_{jm}$$\in\{[a,a+p_{j}:a\in\mathbb{Z})\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M};$ $(34)$
$x_{jm}^{b}$$\in\{[a,a+p_{j}:a\in\mathbb{Z})\}\cup\{\perp\},$
@@
$y_{b}$$\in\{[a,a+p^{f^{b}}:a\in\mathbb{Z})\}\cup\{\perp\},\nobreak\leavevmode$ $\forall\leavevmode\nobreak\ b\in\mathcal{B};$ $(36)$
$y_{bm}$$\in\{[a,a+p^{f^{b}}:a\in\mathbb{Z})\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ b\in\mathcal{B},m\in\mathcal{M};$ $(37)$
@@
${\mathscr{l}}_{mf}$ $=\sum_{j\in\mathcal{J}_{f}}\text{pulse}\left(x_{jm},s_{j}\right),$ $\forall\leavevmode\nobreak\ m\in\mathcal{M},f\in\mathcal{F}.$ $(41)$
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ m\in\mathcal{M},f\in\mathcal{F},j\in\mathcal{J}_{f_{j}}:$ (50)
alwaysIn $\bigl{(}{\mathscr{l}}_{mf},x_{jm},l_{f},u_{f}\bigr{)}.$
Objective function (4) minimizes the TWCT. Constraints (4) ensure that each job is processed on exactly one machine. This can be seen as an assignment of jobs to machines, but now using the global alternative constraint, which exploits the structure of the problem to help the search process. Similarly, constraints (4) ensure that if a job is processed on a given machine, it is processed on such machine in exactly one batch. This can be seen as another assignment of jobs to batches on specific machines, but now using CP scheduling logic. Constraints (4) guarantee that each batch is processed on exactly one machine. This can be seen as another assignment of batches to machines with global constraints. The previous three constraint groups are completely independent and allow the machine assigned to each job, the batch where it is assigned, and the machine assigned to the batch, to be completely unrelated. Constraints (4) remedy this and link them all together, ensuring that if the interval variable $x^{b}_{jm}$ for the job $j\in\mathcal{J}$ in the batch $b\in\mathcal{B}_{f_{j}}$ on the machine $m\in\mathcal{M}$ is present in the final solution, then it must be the case that all the intervals in $V^{b}_{jm}$ must be present in the solution as well (i.e., the batch interval $y_{b}$, the interval of the batch on the machine $y_{bm}$, and the interval of the job on the machine $x_{jm}$). Furthermore, constraints (4) synchronize the start and end of the interval $x^{b}_{jm}$ with the intervals in $V^{b}_{jm}$, if they are present in the final solution. Constraints (4) ensure that only one batch is processed at a time on each machine. Constraints (4) ensure that jobs of only one family are processed at any time on the machines. Finally, constraints (4) guarantee that the load on the machines is within the required batch size limits of the family being processed.
VI-D Breaking symmetries
@@
presenceOf $\left(y_{b}\right)\Rightarrow\texttt{presenceOf}\left(y_{b-1}\right);$ $(51)$
@@
Constraints (VI-D) ensure that the batches of a family are used incrementally, i.e., if a batch $b$ is used, is because all the previous batches have already been used. Constraints (VI-D) go further by ensuring that the start times of the batches also increase with the batches. They use the global constraint startBeforeStart, which ensures that the start time of the first interval variable provided is before the start time of the second interval provided.
@@ VII Computational experiments and results
Table II: The Structure of the instances.
Category Levels Definitions
Jobs 4 $|\mathcal{J}|\in\set{15,25,50,100}$
Families 2 ||$|\mathcal{F}|\in\set{3,5}$
Machines 2 ||$|\mathcal{M}|\in\set{2,3}$
Capacity 1 $u_{f}=50,l_{f}=1,\forall\leavevmode\nobreak\ f\in\mathcal{F}$
Processing times 2 $p^{f}\sim U\set{[\varsigma]},\forall\leavevmode\nobreak\ \varsigma\in\set{5,10}$
Job sizes 2 $s_{j}\sim U\set{[\varsigma]},\forall\leavevmode\nobreak\ \varsigma\in\set{25,50}$
Job weights 2 $\omega_{j}\sim U\set{[\varsigma]},\forall\leavevmode\nobreak\ \varsigma\in\set{5,10}$
Release time multipliers 2 $\lambda\in\set{0.5,1}$
# of instance classes 256
# of instances per class 10
Total # of instances 2,560
The computational experiments consider 2,560 instances following the same process used by Cakici et al. [6] and Ham et al. [18]. Table II presents the structure of the instances. The number of jobs in each instance can be one of four possible values: 15, 25, 50, or 100. The number of families in each instance can be one of two possible values: 3 or 5. The number of machines can be one of two possible values: 2 or 3. The size of the batches of each family must be between 1 and 50 in every instance. The processing times, job sizes, and job weights in each instance follow the distribution $U\set{[\varsigma]}$, which is the discrete uniform distribution over the values of the discrete set $[\varsigma]=\set{1,2,3,\ldots,\varsigma}$, where $\varsigma$ can be either 5 or 10 for the processing times and job weights, or either 25 or 50 for the job sizes. To generate the release times of the jobs, a lower bound of the overall makespan is computed as $C_{max,lb}=\left(\sum_{f\in\mathcal{F}}p^{f}\cdot\lceil N_{f}/u_{f}\rceil\right)/{|\mathcal{M}|}$. Then, this $C_{max,lb}$ is multiplied by $\lambda$ to define the possible set of release times. Hence $r_{j}\sim U\set{[\lambda\cdot C_{max,lb}]}$, where the $\lambda$ multiplier can take one of two possible values: 0.5 or 1. Overall, a total of $256=4\times 2^{6}$ instance classes result from the combination of the values and distributions described above. For each instance class (i.e., combination), 10 random instances were generated with batch size limits from 1 to 50, resulting in a grand total of $2,560$ instances.
The results only compare models that produce feasible solutions. Hence, each one of the generated instances was solved using the following six models:
  • NS+FR: Non-synchronized model by Ham et al. [18] + Feasibility Restoration step described in §V-B.
  • AS: Assign-and-Schedule model inspired by Nascimento et al. [8, 9].
  • AS+SB: AS model + Symmetry-breaking constraints.
  • S: Synchronized model inspired by Cakici et al. [10].
  • RS: The novel Redundant Synchronized model.
  • RS+SB: RS model + Symmetry-breaking constraints.
All the models and FR steps were implemented in Java 8. All the CP models were solved with IBM ILOG CP Optimizer [37] from CPLEX 22.1.1, using its Java interface [38]. All the experiments were run on the PACE Phoenix cluster [39], using machines that run Red Hat Enterprise Linux Server release 7.9 (Maipo) with dual Intel®  Xeon®  Gold 6226 CPU @ 2.70GHz processors, with 24 cores, and 48 GB RAM. Each run uses 8 cores and 16 GB RAM, and is given a time limit of 10 minutes (600 seconds) because a diffusion scheduling system in the semiconductor industry is expected to generate a Gantt-chart schedule every few minutes [18].
Figures 4-6 summarize the results. They group the 2,560 instances in 16 instance classes of 160 instances each, according to their number of jobs, families, and machines. Figure 4 presents the percentage of instances solved to optimality by each model in the top part of the chart, and the percentage of instances where each model could not find a solution in the bottom part. The key takeaway is that the RS model dominates all the other models by solving to optimality the most amount of instances with 15 and 25 jobs, and finding solutions to all the instances with 50 and 100 jobs. No model is able to prove optimality in the larger instances with 50 and 100 jobs. Furthermore, the AS and AS+SB models cannot even find solutions to some of these larger instances.
Figure 4: Percentage of instances solved to optimality (top) and with no solution found (bottom).
Figure 5 presents the average solve times obtained by each model. The stripes around the lines indicate the 95% confidence intervals. On the large instances with 50 and 100 jobs, most of the models reach the time limit given, which explains why none proves optimality in Figure 4. The NS+FR step is the fastest of all models, since the NS model produces infeasible solutions fast, and the FR step recovers their feasibility quickly. Nonetheless, as seen in Figure 4, this approach does not solve to optimality many instances compared to our RS model, which is the second fastest model.
Figure 5: Average solve time (s) by model with 95% confidence intervals.
To gain more insight on the quality of each model, the minimum TWCT is obtained across all models and is denoted by $\text{TWCT}_{i}^{*}$. Then, for each model, the gap on the instance as $|\text{TWCT}_{i,model}-\text{TWCT}_{i}^{*}|/|\text{TWCT}_{i,model}|$ is computed for each model. Figure 6 shows the average gap of each model on the instances where all the models found a solution. The stripes around the lines are the 95% confidence intervals. Although the NS+FR model is fast, its solutions reach up to 10% gap in the larger instances. It is particularly bad in instances with more machines since the FR step recovers the feasibility independently on each machine, being myopic to possible improvements by swapping jobs to other machines. The AS and AS+SB models perform better than the NS+FR model in the smaller instances but slightly worse in the larger ones. This is due to the lack of global constraints that help the search process. The S model performs slightly better than the AS model in the larger instances due to the replacement of the binary variables with interval variables and the addition of more global constraints. From the graph, it becomes clear that the RS model dominates all the other models, providing the best solutions overall. This is due to the redundancy in the structure of the problem and the exclusive use of global constraints. Interestingly, the SB constraints deteriorate the quality of the solutions obtained by the AS and RS models.
Figure 6: Average gap by model with 95% confidence intervals.

### v3

@@ VI-D The Redundant Synchronized Model
$x_{j}$$\in\{[a,a+{\delta}_{j}:r_{j}\leq a\in\mathcal{H})\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J};$ $(47)$
$x_{jm}$$\in\{[a,a+{\delta}_{j}:a\in\mathcal{H})\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ j\in\mathcal{J},m\in\mathcal{M};$ $(48)$
$x_{jm}^{b}$$\in\{[a,a+{\delta}_{j}:a\in\mathcal{H})\}\cup\{\perp\},$
@@
$y_{b}$$\in\{[a,a+{\delta}^{f^{b}}:a\in\mathcal{H})\}\cup\{\perp\},\nobreak\leavevmode$ $\forall\leavevmode\nobreak\ b\in\mathcal{B};$ $(50)$
$y_{bm}$$\in\{[a,a+{\delta}^{f^{b}}:a\in\mathcal{H})\}\cup\{\perp\},$ $\forall\leavevmode\nobreak\ b\in\mathcal{B},m\in\mathcal{M};$ $(51)$
@@
${\mathscr{l}}_{mf}$ $=\sum_{j\in\mathcal{J}_{f}}\text{pulse}\left(x_{jm},{n}_{j}\right),$ $\forall\leavevmode\nobreak\ m\in\mathcal{M},f\in\mathcal{F}.$ $(55)$
@@
$\forall\leavevmode\nobreak\ m\in\mathcal{M},f\in\mathcal{F}:\quad{\mathscr{l}}_{mf}\leq u_{f}.$ $(64)$
Equation (5) defines an interval variable $x_{j}$ of size ${\delta}_{j}$ that represents job $j\in\mathcal{J}$. Equation (5) defines an optional interval variable $x_{jm}$ that represents the option of processing job $j$ on machine $m\in\mathcal{M}$. Equation (5) defines an optional interval variable $x_{jm}^{b}$, also of size ${\delta}_{j}$, that represents the option of job $j$ being batched in batch $b\in\mathcal{B}_{f_{j}}$ and processed on machine $m$. Equation (5) defines an optional interval variable $y_{b}$ of size ${\delta}^{f^{b}}$ that represents batch $b\in\mathcal{B}$. Equation (5) defines an optional interval variable $y_{bm}$ of size ${\delta}^{f^{b}}$ that represents the option of batch $b$ being processed on machine $m$. Equation (5) defines an interval sequence variable $\varphi_{m}$ that represents the sequence of batches scheduled on machine $m$. Equation (5) defines the set $V_{jm}^{b}$ of interval variables that have to be synchronized with interval $x_{jm}^{b}$, if it is present. Equation (5) defines the unique state ${\mathfrak{f}}_{m}$ of the machine $m$ at every point in time. Equation (5) defines a cumulative function ${\mathscr{l}}_{mf}$ that represents the load on machine $m$ with jobs of family $f$.
Objective function (5) minimizes the TWCT. Constraints (5) ensure that each job is processed on exactly one machine. This can be seen as an assignment of jobs to machines, but now using the global alternative constraint, which exploits the structure of the problem to help the search process. Similarly, constraints (5) ensure that if a job is processed on a given machine, it is processed on such machine in exactly one batch. This can be seen as another assignment of jobs to batches on specific machines, but now using CP scheduling logic. Constraints (5) guarantee that each batch is processed on exactly one machine. This can be seen as another assignment of batches to machines with global constraints. The previous three constraint groups are completely independent and allow the machine assigned to each job, the batch where it is assigned, and the machine assigned to the batch, to be completely unrelated. Constraints (5) remedy this and link them all together, ensuring that if the interval variable $x^{b}_{jm}$ for the job $j\in\mathcal{J}$ in the batch $b\in\mathcal{B}_{f_{j}}$ on the machine $m\in\mathcal{M}$ is present in the final solution, then it must be the case that all the intervals in $V^{b}_{jm}$ must be present in the solution as well (i.e., the batch interval $y_{b}$, the interval of the batch on the machine $y_{bm}$, and the interval of the job on the machine $x_{jm}$). Furthermore, constraints (5) synchronize the start and end of the interval $x^{b}_{jm}$ with the intervals in $V^{b}_{jm}$, if they are present in the final solution. Constraints (5) ensure that only one batch is processed at a time on each machine. Constraints (5) ensure that jobs of only one family are processed at any time on the machines. This global constraint is only available in CP Optimizer. When implementing in other solvers, removing this constraint does not impact the feasibility of the solution achieved since constraints (5) already ensure this as well. Finally, constraints (5) guarantee that the load on the machines is within the required batch size limits of the family being processed.
To minimize the makespan, objective function (5) should be replaced by objective function (9), which minimizes the maximum end time of the jobs.
VI-E Breaking symmetries
@@
presenceOf $\left(y_{b}\right)\leq\texttt{presenceOf}\left(y_{b-1}\right);$ $(65)$
@@
Constraints (VI-E) ensure that the batches of a family are used incrementally, i.e., if a batch $b$ is used, is because all the previous batches have already been used. Constraints (VI-E) go further by ensuring that the start times of the batches also increase with the batches. They use the global constraint startBeforeStart, which ensures that the start time of the first interval variable provided is before the start time of the second interval provided. This constraint can be implemented in other solvers using reification.
@@ VII Computational experiments and results
(a) % of instances without solution
(b) Average gap (%) with 95% CI
Figure 6: Results minimizing the TWCT with CP Optimizer
(a) % of instances without solution
(b) Average gap (%) with 95% CI
Figure 7: Results minimizing the TWCT with CP-SAT
(a) % of instances solved to optimality
(b) Average solve time (s)
(c) % of instances without solution
(d) Average gap (%) with 95% CI
Figure 8: Results minimizing the $C_{\max}$ with CP Optimizer.
(a) % of instances solved to optimality
(b) Average solve time (s)
(c) % of instances without solution
(d) Average gap (%) with 95% CI
Figure 9: Results minimizing the $C_{\max}$ with CP-SAT.
The computational experiments consider 1,000 instances generated following a similar process to that of Cakici et al. [6] and Ham et al. [18]. The number of jobs in each instance is one of four possible values: 50, 100, 150, or 200. Likewise, the number of families and the number of machines in each instance can be one of five possible values: 4, 5, 6, 8, or 10. Overall, there are ${100}=4\times 5\times 5$ instance classes, and for each class, 10 random instances were generated, totaling 1,000 instances.
For every instance, the processing time for each family $f\in\mathcal{F}$ and the weight for each job $j\in\mathcal{J}$ follow the distribution ${\delta}^{f},\omega_{j}\sim U\set{[{10}]}$, where $U\set{[\varsigma]}$ denotes the discrete uniform distribution over the set $[\varsigma]=\{1,2,\dots,\varsigma\}$. Job sizes are drawn from ${n}_{j}\sim U\set{[25]}$. Release times are generated by first computing a lower bound of the overall makespan as
$C_{\max,lb}=\frac{\sum_{f\in\mathcal{F}}{\delta}^{f}\cdot\lceil N_{f}/u_{f}\rceil}{|\mathcal{M}|},$
and then setting $r_{j}\sim U\set{[C_{\max,lb}]}$. Finally, the maximum batch size for each family is generated as
$u_{f}\sim U(\{10,15,20,25,30,35,40,45,50,75,100,125,150\}).$
Models and Solvers
Each instance is solved by minimizing two objectives independently—TWCT and $C_{\max}$—using the following seven models:
  • A: Aligned model by Ham et al. [18].
  • AU: Automaton model from Section VI-A.
  • AS: Assign-and-Schedule model from Section VI-B.
  • AS+SB: AS model with symmetry-breaking (SB) constraints.
  • S: Synchronized model from Section VI-C.
  • RS: Redundant Synchronized model from Section VI-D.
  • RS+SB: RS model with symmetry-breaking constraints.
To showcase the flexibility of our modeling approach, we used two solvers:
  • Commercial solver: IBM ILOG CP Optimizer [39] from CPLEX v.22.1.1,
  • Open-source solver: Google OR-Tools CP-SAT [43, 41] v.9.11.
Table II summarizes the models implemented in each solver. The A model was only implemented in CP Optimizer because its state functions, alwaysEqual global constraint, and alignment parameters are only available int. On the other hand, the AU model was only implemented in CP-SAT because the automaton global constraint required is not available in CP Optimizer (although it can be implemented in Choco-solver [22] and MiniZinc [42]). The AS, S, and RS models (and their variants with SB constraints) are implemented in both solvers.
Table II: Models implemented by solver
Solver Models Total
CP Optimizer  A , AS, AS+SB, S, RS, RS+SB 6
CP-SAT AU, AS, AS+SB, S, RS, RS+SB 6
[…]
Figures 69 summarize our results. The graphs for TWCT focus on Figures 6 and 7, while Figures 8 and 9 detail the results for the $C_{\max}$ objective. Each graph presents the following characteristics:
  • Instance Grouping: the stacked line charts group instances by the number of jobs (indicated on the right-hand side) and by the combination of families and machines along the horizontal axis.
  • Data Aggregation: each marker represents the aggregated result (over 10 instances) of a given instance class.
  • Model Comparison: each series corresponds to one of the six models implemented in each solver, yielding metrics from 6,000 experiments per graph.
The metrics reported for each model include:
  1. 1.
    Percentage of instances solved to optimality.
  2. 2.
    Percentage of instances where no solution was found
  3. 3.
    Average gap (%) with its 95% confidence interval (CI).
  4. 4.
    Average solve time (s) with its 95% CI.
The gap for each instance is computed as follows. First, we obtain the minimum objective value $z_{i,obj}^{*}$ across all models and both solvers for each instance $i$ and objective $obj\in\{TWCT,C_{\max}\}$. Then, for a given model and solver, the gap is defined as $\frac{|z_{i,obj,\text{model, solver}}-z_{i,obj}^{*}|}{|z_{i,obj}^{*}|}$.
VII-A Results with the TWCT objective
Figures 6 and 7 present the results minimizing the TWCT with CP Optimizer and CP-SAT, respectively. Under the TWCT objective, all the models reached the time limit of 1,200 seconds. Hence, the graphs that show the average solve times of the models for both solvers are omitted. Additionally, no model declared optimality on any instance, with the exception of the AS+SB model implemented in CP-SAT, which declared optimality in only one instance with 50 jobs, 4 families, and 10 machines. For this reason, the graphs that show the percentage of instances solved to optimality for both solvers are also omitted.
Figure 6 focuses on the results obtained with CP Optimizer. In particular, Figure 6(a) reports the percentage of instances for which each model failed to find a solution. For instances with 50 and 100 jobs, all models successfully found solutions. However, as the number of jobs increases, the failure rate rises notably for the AS, AS+SB, and RS+SB models. The AS model (shown in dark blue) exhibits a distinct seasonal pattern: its performance deteriorates as the number of machines increases, peaking at 10 machines—for instance, no solutions were found for any instance with 200 jobs, 6 families, and 10 machines. The AS+SB model shows a similar seasonal behavior around 10 machines but also fails to find solutions in instances with fewer machines. In contrast, the RS+SB model follows this seasonal behavior yet performs better in larger instances, likely due to randomness in the instance generation. Notably, the A, S, and RS models were able to find solutions in all instances.
Figure 6(b) presents the average gap (%) along with its 95% confidence interval for each model in CP Optimizer. Although the A model did not prove optimality within the time limit in any instance, it consistently finds the best solutions, evident by the constant 0% gap of the black series. In contrast, the AS and AS+SB models exhibit the worst average gaps, with wider confidence intervals. This is partly due to their failure to find solutions in several instances, resulting in fewer data points for averaging. Notably, a series break for these models appears at 200 jobs, 6 families, and 10 machines, reflecting a complete failure to solve any instance in that class. Furthermore, the consistently lower gaps of the AS+SB model compared to the AS model underscore the beneficial impact of incorporating SB constraints in the search process.
The AS and AS+SB models perform worse in CP Optimizer because of their structure. Although the beauty of these models is their clear distinction between assignment decisions with binary variables and scheduling decisions with interval variables and global constraints, this separation reduces the ability of CP Optimizer to prune the domains of the scheduling variables. The noOverlap constraints, in charge of handling the scheduling logic, provide the CP engine with a global view of the disjunctive constraints and effectively reduce the search space. However, constraints (3)-(3), in charge of handling the assignment while respecting the batch limits, are local in nature, reduce the overall propagation, and provide weak information to the global scheduling constraints. As a result, this myopic structure heavily impacts these models’ performance. Instead, the S and RS models, which replace the binary assignment with interval variables and global constraints, provide better information to the search process, allowing them to find solutions to all the instances and find solutions that are close to be the best.
The extra SB constraints of the RS+SB impact this model’s ability to find solutions in larger instances because they become harder to satisfy within the time limit. Note that the SB constraints impose additional requirements on the identifier of the batches. This is only useful for constraints (5) to determine which batches to use. However, the implication constraints (5) only enforce the presence of a batch if the interval of any job in such batch on any machine is also present, not vice-versa. Hence, the SB constraints don’t really provide better information to the CP engine during the search process to determine which job intervals to choose, precisely because of this directionality of the implications. This in turn makes the SB constraints harder to satisfy, causing more failures during the search.
Figure 7 focus on the results of CP-SAT, which combines CP methods and satisfiability (SAT) methods in the search process. SAT methods benefit from binary structures in the problem. For this reason, in this solver, the AS model is able to find solutions on all the instances, as evidenced in Figure 7(a). The distinction between assignment and scheduling decisions of the AS model provides the proper structure for both CP and SAT methods to properly work, allowing the AS model to find solutions in all the instances; and even declare optimality in one of them. In contrast, the RS+SB model is unable to find solutions in most of the largest instances, precisely because of the additional SB constraints that become harder to satisfy within the time limit. The RS model is unable to find solutions in the largest instances with 200 job, precisely because of the amount of job-in-batch decisions that need to be made. On the other hand, the AU model clearly shows seasonal behavior around 10 machines.
When examining the average gaps in CP-SAT from Figure 7(b), it becomes evident that no model consistently matches the best results obtained by the A model in CP Optimizer. Nonetheless, the gaps are very close. The AU gap increases with the size of the instances. This is because of the $\tau_{mt}$ variables that represent the transition values provided to the automaton at every time step of the scheduling horizon, which increases with the number of jobs. Additionally, the element constraints set the appropriate transition values at the start times of the jobs on each machine. However, they do so only if the jobs are actually processed on such machine. While this could be seen as a reified constraint, it is not. Instead, this is indirectly achieved with the modified start times $\hat{s}_{jm}$ and modified transition values $\hat{\tau}_{jm}$. When the instances grow, more of these modified decisions have to be made, negatively impacting the performance of these element constraints, and causing the average gap increase.
In CP-SAT, it becomes evident that satisfying the SB constraints within the time limit is difficult for both the AS and RS models. The SB constraints not only impose additional relationships between the boolean presence of the batches, but also impose additional relationships between their start times, which are only enforced if the batches are actually present. This is achieved through reification. Thus, these SB constraints modify the nice boolean structure that makes the SAT methods in CP-SAT shine, exchanging it with reified constraints that negatively impact the models’ performance. This behavior has been documented in the CP-SAT Primer by Krupke et al. [46].
The best gaps in CP-SAT belong to the AS model, followed by the RS model, very close to 0%. Interestingly, the very same reason that makes The AS model model fall back in CP Optimizer, makes it shine in CP-SAT instead. The clear distinction between assignment and scheduling logic is well handled in CP-SAT thanks to the SAT component of this solver. Additionally, the RS model consistently produces comparable gaps of the AS model. However, as seen in Figure 7(a), the RS model is not able to find solutions in all the instances.
VII-B Results with the makespan objective
Figures 8 and 9 present the results for minimizing $C_{\max}$ using CP Optimizer and CP-SAT, respectively. The results clearly indicate that the $C_{\max}$ objective is easier to solve than TWCT, primarily because it is insensitive to the processing order of jobs—only the maximum completion time matters. As a consequence, several models are able to prove optimality in some instances, as shown in Figures 8(a) and 9(a), and they achieve these optimal solutions before the time limit is reached, as evidenced by the solve times in Figures 8(b) and 9(b). Interestingly, the seasonal behavior observed in these figures suggests that instances with a higher number of machines are easier to solve, since fewer jobs per machine simplify the identification of the maximum completion time. In CP Optimizer, Figures 8(a) and 8(b) further reveal that the A model scales exceptionally well, finding optimal solutions in a larger fraction of instances and in less time—thanks largely to the effective domain pruning provided by the alwaysEqual global constraint with alignment parameters.
Figures 8(c) and 9(c) (percentage of instances without a solution) and Figures 8(d) and 9(d) (average gap with 95% CI) tell a story similar to that of the TWCT objective. In CP Optimizer, the clear separation between assignment and scheduling decisions in the AS and AS+SB models hampers performance, whereas this separation benefits CP-SAT. In both solvers, the RS+SB model struggles, as the additional symmetry-breaking constraints become increasingly difficult to satisfy. The AS+SB model exhibits a similar issue: although the SB constraints help prove optimality for small instances in CP Optimizer, they become counterproductive as the instance sizes grow, leading to poorer gap values. In CP-SAT, the necessary reification in these constraints generally degrades performance.
The AU model performs well for instances with up to 150 jobs; however, when scaling to 200 jobs, it faces challenges in finding solutions for instances with 8 or 10 machines. Finally, the RS and S models show average performance: they do not match the AS model in CP-SAT, yet they consistently outperform the RS+SB model.